OpenStudy (anonymous):

What is the solution to the following system of equations? y = x2 + x + 8 y = x2 + 2x + 5 last one

OpenStudy (anonymous):

@ivettef365

HanAkoSolo (jamierox4ev3r):

is the 2 in x2 an exponent (x^2)?

OpenStudy (anonymous):

yes

HanAkoSolo (jamierox4ev3r):

x^2+x+8=x^2+2x+5

HanAkoSolo (jamierox4ev3r):

because both are equal to "y", they are equal to each other

OpenStudy (anonymous):

5,8

HanAkoSolo (jamierox4ev3r):

x^2+x+8=x^2+2x+5 -x^2 -x^2 x+8=2x+5 -x -x 8=x+5 -5 -5 x=3

HanAkoSolo (jamierox4ev3r):

x=3. where did you get 5,and 8?

OpenStudy (anonymous):

3,20 ?

OpenStudy (anonymous):

didnt mean to say that

HanAkoSolo (jamierox4ev3r):

lol okay

HanAkoSolo (jamierox4ev3r):

then you plug in the x variable to both equations idk if it is 20 but let me check for you

HanAkoSolo (jamierox4ev3r):

y=3^2+3+8 y=9+3+8 y=20

HanAkoSolo (jamierox4ev3r):

first one seems good, let me check the next

OpenStudy (anonymous):

thank you :)

HanAkoSolo (jamierox4ev3r):

y=3^2+2(3)+5 y=9+6+5 y=20

HanAkoSolo (jamierox4ev3r):

so your answer is (3,20) you were correct good job :D